y= x-4
y= -x+2 how to solve by substition

Answer by rapaljer(4671)

Substitute one of the y's in place of the other one.
Instead of writing
y=-x+2

replace this y with x-4:
y=-x+2
x-4=-x+2

Add +x to each side:
x+x-4=-x+x+2
2x-4=2

Add +4 to each side:
2x-4+4=2+4
2x=6
Divide both sides by 2
x=3

Now, you can find y in either equation. Take the first one:
y=x-4
y=3-4
y=-1

Check by substituting both values into the other equation
y=-x+2
-1=-3+2 It checks!!

For additional help with Systems of Equations, please see my own website by clicking on my tutor name "rapaljer" anywhere in algebra.com. Then click on my website. On my Homepage, look for the link "Basic, Intermediate, and College Algebra: One Step at a Time." Then choose "Basic Algebra", then "Chapter 4" and look for "Section 4.07 Systems of Equations. I have a complete explanation of Logarithms that I think you will find my explanation a lot easier to understand than traditional textbooks. Besides this, many of my exercises are solved IN COLOR in the MATH IN LIVING COLOR pages that go with this.

I have many other topics from Basic, Intermediate, and College Algebra that are explained in detail, with many of these topics explained on video as well. Everything on the website is FREE.

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus